OpenStudy (anonymous):

how do I find the x intercepts of a parabola given by this equation; y = 2xsquared-4x

OpenStudy (phi):

you find the x's where y=0 in other words, solve \[ 2x^2 -4x = 0 \] I would first divide both sides by 2 then factor x from the two terms on the left

OpenStudy (anonymous):

would you do a step by step for I am not understanding the example in my book

OpenStudy (phi):

can you divide both sides by 2 ? what do you get ?

OpenStudy (phi):

yes x^2 -2x = 0 next factor x from each term to get x(x-2)=0

OpenStudy (phi):

now use this (nifty !) idea: if A*B =0 then either A=0 or B=0 (or both at the same time are zero) in other words if you have x * (x-2) = 0 then it could be x=0 or (x-2)=0 what two values make x(x-2) = 0 ? clearly x=0 is one what makes x-2 = 0 true ?